OpenStudy (anonymous):

Find an equation line of the tangent line to the given curve at the given point.

OpenStudy (anonymous):

\[y=sinx \] at (pi/6 , 1)

OpenStudy (anonymous):

y-1= cosx(x-(Pi/6))

OpenStudy (anonymous):

find the slope of tangent first y=sinx diff it once dy/dx =cosx put x=pi/6 dy/dx = cos pi/6 =sqrt3 /2 tangent: y=sqrt3 /2 x +c put x =pi/6 and y=1 1=(sqrt3 /2)(pi/6)+c you can then find c put the value you get into y=sqrt3 /2 x +c We can then get the equation of tangent

OpenStudy (anonymous):

Answer is : \[6\sqrt{3x} - 6y - \pi \sqrt{3} + 6 = 0\]